OpenStudy (anonymous):

y = 1/2x + 1 and y= -1/2x + 3

OpenStudy (anonymous):

1/2x+1=-1/2x+3 2/2x=3-1 x=2

OpenStudy (anonymous):

I dont this that right im doing system graphing and I need to find the x and the y for where they cross